Haloacetic acids are by-products of drinking water disinfection. In Superior CharterTownship, the level of haloacetic acids present in the water was 13 μg/L. Assume haloacetic acids are monoprotic with an average molar mass of 150 g/mol, and the only source of H+ in the water sample. What volume of 0.5 M calcium hydroxide would be required to neutralize a water sample of 24.5 mL?
Answer #1

Haloacetic molarity is calculated:

M = 13 ug / L * (1 g / 1x10 ^ 6 ug) * (1 mol / 150 g) = 8.7x10 ^ -8 M

The volume of Ca (OH) 2 required is calculated:

V Ca (OH) 2 = Ca * Va / 2 * Cb = 8.7x10 ^ -8 M * 24.5 mL / 2 * 0.5 M = 2.13x10 ^ -6 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.